# Berkeley, CA 94720, USA

### v2

@@ 5.2 Examples
In the examples below, we will take ${\mathfrak{n}}$ string sources each wound once along the $1$-cycle of $S^{3}/\Gamma$ associated to the generator $r^{2}\in\text{Dic}_{k}$. (We chose the element $r^{2}$ so as to get nontrivial results.) From (80) we have
$\text{Index}_{k,{\mathfrak{n}}}$ $=\sum_{\begin{array}[]{c}\text{irrep $R_{m}$}\\ \dim R_{m}=2\\ \end{array}}\frac{1}{2}\tau(\Gamma,R_{2m}\oplus{\mathbf{1}_{{{2}}}})\chi_{R}(r^{2})^{{\mathfrak{n}}}$ (86)
$=\sum_{m=1,3,5,\dots}\frac{1}{2}\tau(\text{Dic}_{k},{\mathbf{2}_{{2m}}}\oplus{\mathbf{1}_{{{2}}}})\left[2\cos\frac{2\pi m}{k}\right]^{\mathfrak{n}},$ (87)
and $m$ goes up to $k-1$ for even $k$ and up to $k-2$ for odd $k$. Here we can use (70) and set
Thus,
$\text{Index}_{k,{\mathfrak{n}}}=\frac{2^{3+{\mathfrak{n}}}}{k}\sum_{m=1,3,5,\dots}\cos^{\mathfrak{n}}\left(\frac{2\pi m}{k}\right)\sin^{2}\left(\frac{\pi m}{k}\right).$ (89)
@@
$\text{Index}_{2,{\mathfrak{n}}}=\frac{1}{2}\tau(\text{Dic}_{2},{\mathbf{2}_{{2}}}\oplus{\mathbf{1}_{{{2}}}})\left[2\cos\pi\right]^{\mathfrak{n}}\\ =\begin{cases}4,&\quad\text{for ${\mathfrak{n}}=0$},\\ (-2)^{{\mathfrak{n}}}4,&\quad\text{otherwise}\end{cases}$
@@
$\text{Index}_{3,{\mathfrak{n}}}=\frac{1}{2}\tau(\text{Dic}_{3},{\mathbf{2}_{{2}}}\oplus{\mathbf{1}_{{{2}}}})\left[2\cos\tfrac{2\pi}{3}\right]^{\mathfrak{n}}=(-1)^{{\mathfrak{n}}}2$
@@
$=\frac{1}{2^{{\mathfrak{n}}-1}}\left([\sqrt{5}-1]^{{\mathfrak{n}}}+[-1-\sqrt{5}]^{{\mathfrak{n}}}\right)$
$\longrightarrow-2,4,-6,10,-16,26,\ldots,\quad\text{for ${\mathfrak{n}}\geq 1$}$
@@
suggests a fusion relation whereby two strings can fuse to a linear combination of one or zero strings. We derive these recurrence relations for general $k$ below.

### v3

@@ 5.2 Examples
In the examples below, we will take $\mathfrak{n}$} string sources each wound once along the $1$-cycle of $S^{3}/\Gamma$ associated to the generator $r^{2}\in\text{Dic}_{k}$. (We chose the element $r^{2}$ so as to get nontrivial results.) We also write $\text{Index}_{k,{\mathfrak{n}}}$, short for $\text{Index}_{k,{\mathfrak{n}}}(p_{1},\cdots,p_{\mathfrak{n}})$ in (88). Thus,
$\text{Index}_{k,{\mathfrak{n}}}=\frac{2^{{\mathfrak{n}}+3}}{k}\sum_{m=1,3,5,\dots}\cos^{\mathfrak{n}}\left(\frac{2\pi m}{k}\right)\sin^{2}\left(\frac{\pi m}{k}\right),$ (92)
and $m$ goes up to k1k-1 for even $k$ and up to k2k-2 for odd $k$. Define
[…]
With basic algebraic manipulations such as a binomial expansion and summation of a geometric series, we can rewrite the index as
$\text{Index}_{k,{\mathfrak{n}}}$ $=$ $2\sum_{l=0}^{{\mathfrak{n}}}\binom{{\mathfrak{n}}}{l}(-1)^{2({\mathfrak{n}}-2l)/k}\delta_{k}(2({\mathfrak{n}}-2l))$ (94)
$-\sum_{l=0}^{{\mathfrak{n}}+1}\binom{{\mathfrak{n}}+1}{l}(-1)^{2({\mathfrak{n}}+1-2l)/k}\delta_{k}(2({\mathfrak{n}}+1-2l)),$
which for odd $k$ simplifies to
$\text{Index}_{k,{\mathfrak{n}}}=2\sum_{l=0}^{{\mathfrak{n}}}\binom{{\mathfrak{n}}}{l}\delta_{k}({\mathfrak{n}}-2l)-\sum_{l=0}^{{\mathfrak{n}}+1}\binom{{\mathfrak{n}}+1}{l}\delta_{k}({\mathfrak{n}}+1-2l)\qquad(k\in 2\mathbb{Z}+1),$ (95)
The expression for the index in (94) is manifestly an integer. We will present additional manifestly integral formulas for specific values of $2\leq k\leq 6$ below.
@@
$\text{Index}_{2,{\mathfrak{n}}}=\frac{1}{2}\tau(\text{Dic}_{2},{\mathbf{2}_{{2}}}\oplus{\mathbf{1}_{{{2}}}})\left[2\cos\pi\right]^{\mathfrak{n}}\\ =(-2)^{{\mathfrak{n}}}4,$
for all $\mathfrak{n}$}.
@@
$\text{Index}_{3,{\mathfrak{n}}}=\frac{1}{2}\tau(\text{Dic}_{3},{\mathbf{2}_{{2}}}\oplus{\mathbf{1}_{{{2}}}})\left[2\cos\tfrac{2\pi}{3}\right]^{\mathfrak{n}}=(-1)^{{\mathfrak{n}}}2,$
@@
$=\frac{1}{2^{{\mathfrak{n}}+2}\sqrt{5}}\left((\sqrt{5}-1)^{n+1}-(-1-\sqrt{5})^{n+1}\right)$
$\longrightarrow 2,-2,4,-6,10,-16,26,\ldots,\quad\text{for ${\mathfrak{n}}\geq 0$}$
@@
suggests a fusion relation whereby two strings can fuse to a linear combination of one or zero strings. We derive these recurrence relations for general $k$ in §5.3 below.